# Krylov Complexity and Dynamical Phase Transition in the quenched LMG model

### v1

@@ Abstract
Investigating the time evolution of complexity in quantum systems entails evaluating the spreading of the system’s state across a defined basis in its corresponding Hilbert space. Recently, the Krylov basis has been identified as the one that minimizes this spreading. In this study, we develop a numerical exploration of the Krylov complexity in quantum states following a quench in the Lipkin-Meshkov-Glick model. Our results reveal that the long-term averaged Krylov complexity acts as an order parameter when the quench originates from a zero magnetic field. It effectively discriminates between the two dynamic phases induced by the quench, sharing a critical point with the conventional order parameter. Additionally, we examine the inverse participation ratio and Shannon entropy in both the Krylov basis and the energy basis. A matching dynamic behavior is observed in both bases when the initial state possesses a specific symmetry. This behavior is analytically elucidated by establishing the equivalence between the Krylov basis and the pre-quench energy eigenbasis.
@@ I Introduction
Quantifying the complexity of physical phenomena has been a long-standing area of interest [1]. Various measures, such as the entropy generated during the time evolution of a quantum state [2], algorithmic randomness [3], and quantum Kolmogorov complexity [4], have been employed to characterize complexity. While many existing complexity measures can capture certain features, some fail in encompassing other aspects or even exhibit ambiguity due to the specific basis choices involved in their definition.
@@
There is a widely accepted notion that chaotic systems exhibit greater complexity compared to integrable ones, attributed in part to phenomena such as the butterfly effect [5, 6]. In the realm of quantum systems, this effect is typically characterized by the exponential growth of the noncommutativity of local operators over time, as quantified by out-of-time order correlators. However, this exponential increase is not consistently observed in chaotic systems and may even manifest in nonchaotic ones [7, 8, 9], rendering this feature elusive.
@@
A novel measure of complexity has recently been proposed, relying on the spread of operators across a specifically ordered basis known as the Krylov basis [10]. This basis provides the visualization of dynamics as a single particle moving in a semi-infinite chain, with hopping terms corresponding to Lanczos coefficients obtained during the computation of the Krylov basis [11]. A large body of literature has followed, addressing fundamental properties of this measure [12, 13, 14, 15, 16, 17, 18, 19, 20], its generalizations for other kinds of evolution [21, 22, 23, 24, 25, 26], and applications, e.g., to characterize phases of matter [27, 28] and for quantum control [29, 30]. The Krylov complexity involves a choice of inner product between operators, raising questions about the optimal inner product for minimizing complexity. This challenge was addressed with the extension of the Krylov complexity for quantum state evolution presented in Ref. [31]. It was shown that the Krylov basis minimizes the proposed complexity definition. As a result, the Krylov complexity of quantum states, also known as the spread complexity, emerges as an unambiguous measure, eliminating the need for an inner product choice. It has been explored in the SYK model [10, 13], the XXZ model [15], the Su-Schrieffer-Heeger model [28], evolutions governed by dynamical symmetry groups associated with specific Lie algebrass [16, 18, 23], and random matrix models [18, 31], among others. Additionally, it has been applied to analyze the transition from integrability to a chaotic regime [20, 32].
@@
The exploration of dynamics following a global quench in quantum many-body systems is widely used to probe a system in nonequilibrium statistical mechanics [33]. Global quenches induce rapid departures from equilibrium, along with coherence in the energy spectrum, producing intriguing effects, such as dynamical phase transitions that delve into the criticality of quantum many-body system dynamics [34, 35]. Dynamical quantum phase transitions (DQPTs) arise in quenched many-body systems and manifest as cusps in the Loschmidt echo or survival probability, i.e., non-analytic behavior in the time domain for certain initial states. This paper contributes to their study by investigating the evolution of Krylov complexity in quantum states following a quench in the paradigmatic Lipkin-Meshkov-Glick (LMG) model [36]. The LMG model, known to exhibit dynamical phase transitions depending on the quench [37], reveals that the long-time averaged Krylov complexity serves as an order parameter for the dynamical phase transition when the quench starts at null magnetic field. It exhibits distinct behavior over time in both dynamical phases induced by the quench, and remarkably mirrors the oscillations and revivals seen in one of the components of the total magnetization, a conventional dynamical order parameter in spin systems.
@@
In order to understand this fact, we conduct a numerical study of the inverse participation ratio (IPR) and the Shannon entropy in both the Krylov basis and the pre-quench energy basis over time. Remarkably, these two quantities display identical behavior in both bases. Finally, by deriving the Krylov basis for the case of a quench starting at null magnetic field, we analytically demonstrate its equivalence to the pre-quench energy basis. This analytical result explains the identical temporal behavior of IPR and Shannon entropy in these two bases and also why Krylov complexity is an order parameter for the considered model.
@@
The paper is structured as follows: Section II provides a review of key concepts related to dynamical phase transitions while Section III introduces the theory of Krylov complexity of quantum states. The goal of these two sections is twofold: setting the notation of the paper and make it self-contained. Section IV offers an overview of the dynamical phase transition in the LMG model and Section V presents our results. We present our conclusions in Sec. VI.
@@ II Dynamical Quantum Phase Transition
also know as the survial amplitude [38, 39, 40, 41] and the Loschmidt amplitude [42]. The characterization of the DQPT relies on a formal analogy between $\mathcal{G}(t)$ and a special case of the boundary partition function $Z_{B}=\langle\psi_{1}|e^{-RH}|\psi_{2}\rangle$, with $R$ being the distance between the boundaries $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ [43]. Therefore, formally, $\mathcal{G}(t)$ is the boundary partition function with $R=it$ and $|\psi_{1}\rangle=|\psi_{2}\rangle=|\psi_{0}\rangle$. The corresponding probability
@@
The DQPT is then defined in terms of the Fisher, or Lee-Yang zeros [44, 45] of this partition function. Consider the return amplitude with complex time $t\to z=t+i\tau$. A transition occurs every time the zeros of $\mathcal{G}(z)$ crosses the real-time axis of the complex plane, and these crossings indicate the critical times of the dynamics. This analogy with the equilibrium quantum phase transition leads us to another important quantity in the dynamical case, a quantity that can be understood as the dynamical version of the free energy, the rate function
@@
with $N$ being the size of the system. At the critical times, $\mathcal{L}(t)$ is zero and the rate function becomes nonanalytic. It is important to observe that this behaviour only emerges in the thermodynamic limit where $N\rightarrow\infty$ and the Fisher zeros accumulate in a line or in a continuous region [34].
The above-described phenomenon was termed DPT-II to differentiate it from another related phenomenon, known as DPT-I. DPT-II was initially investigated in the transverse field Ising model [42], and since then, it has been explored in various quantum many-body systems. These encompass short to long-range interacting quantum spin chains [46, 47, 37, 48, 49], two-dimensional spin systems [50, 51], nonintegrable systems [52], and optical systems [53], to name a few examples. For an in-depth review of DPT-II, covering both theoretical studies and experimental realizations, see Ref. [34].
@@
While DPT-II is investigated through the lens of the return probability, DPT-I is featured by the Landau order parameter of the system. This order parameter, a quantity whose derivative undergoes nonanalytic changes at a critical point of the quench parameter, serves to distinguish dynamical phases. Typically, the order parameter involves the time average of some physical quantity, such as the magnetization in spin chains. It is noteworthy that this dynamical critical point might not align with the critical point of the equilibrium QPT. DPT-I has been examined in various systems and we point the reader to Refs. [35, 54] for comprehensive reviews on both its theoretical and experimental aspects.
@@
More recently, a connection between DPT-I and DPT-II and the excited-state quantum phase transition (ESQPT) in quantum many-body systems with infinite-range interactions has been established [56]. The authors in Ref. [56] define a generalized microcanonical ensemble by introducing three noncommuting charges, which considers the presence of dynamical order parameters. In the thermodynamic limit, these order parameters are non-zero if the energy $E$ is less than the ESQPT critical energy $E_{c}$, but they always become zero if $E>E_{c}$. Concerning DPT-II, it was demonstrated that non-analyticities in the rate function $r(t)$ only occur if the energy of the system after the quench is greater than $E_{c}$, being prohibited if $E<E_{c}$. The same theoretical framework was applied to a finite-range interacting system in Ref. [57].
@@ III Krylov Complexity of Quantum States
Now we introduce a notion of complexity of the quantum dynamics based on how deep $|\psi_{t}\rangle$ spreads over the system Hilbert space. For this reason, this quantity is also referred as spread complexity [31, 28].
@@
The successive applications of $H$ to the initial state generates the set of quantum states
@@
which describes the spreading of the initial state over the Hilbert space during time evolution. The subspace spanned by $H^{n}|\psi_{0}\rangle\}_{n\geq 0}$ is known as the Krylov space. The measure of complexity that we consider in this work was introduced in Ref. [31] and it is given by
$C_{\mathcal{B}}(t)=\sum_{n}c_{n}|\langle\psi_{t}|B_{n}\rangle|^{2}$ (6)
where $\mathcal{B}$ represents a basis in the Hilbert space, whose elements are denoted as ${|B_{n}\rangle}_{n}$. Intuitively, we expect that complex dynamics lead to larger spreads over this basis. A basis independent complexity measure can be obtained from $C_{\mathcal{B}}(t)$ by performing a special minimization process, resulting in
@@
Of course, it is always possible to construct a basis at the initial time such that $|\psi_{0}\rangle=|B_{0}\rangle$ has a non-zero overlap only with one state of the basis, thereby minimizing complexity. Instead, we consider a functional minimization of (6) which takes into account the spread of the state. For this purpose, it is natural to look at the set $\{H^{n}|\psi_{0}\rangle\}_{n\geq 0}$ spanning the Krylov space, since it includes only the portion of the Hilbert space actually visited by the system over time evolution.
The Krylov complexity, a precedent of the spread complexity, was first introduced in Ref. [10] in the context of operator evolution, i.e., in the Heisenberg picture. Based on this concept, an universal hypothesis for the maximum growth of local operators in quantum many-body systems was presented. Later, Balasubramanian and collaborators [31] extended the idea to the evolution of quantum states, in the Schrödinger picture, and proved that the basis which minimizes (6) for $c_{n}=n$ is the so-called Krylov basis. The Krylov basis is generated performing Gram-Schmidt orthogonalization on the set $\{H^{n}|\psi_{0}\rangle\}_{n\geq 0}$. We denote the Krylov basis as $\mathcal{K}$ and its elements as $\{|K_{n}\rangle\}_{n}$.
Alternatively, the Krylov basis can be generated using the Lanczos algorithm [11, 58], which is a well known recursive method used to generate orthogonal basis. Starting with the initial state as the first Krylov state $|K_{0}\rangle=|\psi_{0}\rangle$, the next state is obtained as $|K_{1}\rangle=\frac{1}{b_{1}}H|K_{0}\rangle$ with $b_{1}=\langle K_{1}|K_{1}\rangle^{1/2}$ being the normalization constant. The subsequent states $\{|K_{n}\rangle\}_{n\geq 2}$ are generated via the following recursion method
$|A_{n}\rangle=H|K_{n-1}\rangle-a_{n}|K_{n-1}\rangle-b_{n-1}|K_{n-2}\rangle$ (8)
@@
Isolating the first term in the right hand side of Eq. (8),
$H|K_{n-1}\rangle=a_{n}|K_{n-1}\rangle+b_{n}|K_{n}\rangle+b_{n-1}|K_{n-2}\rangle,$ (11)
we observe that the hamiltonian is tridiagonal in the Krylov basis $\{|K_{n}\rangle\}_{n}$.
The authors of Ref. [31] proved that the Krylov basis minimizes Eq. (6) in a very specific way. Formally, let
@@
We say that $S_{\mathcal{B}_{1}}<S_{\mathcal{B}_{2}}$ if there is some $k$ such that $C^{(m)}_{\mathcal{B}_{1}}=C^{(m)}_{\mathcal{B}_{2}}$ for $m<k$ and $C^{(m)}_{\mathcal{B}_{1}}<C^{(m)}_{\mathcal{B}_{2}}$ for $m=k$. Thus, for any basis $\mathcal{B}$, $S_{\mathcal{K}}\leq S_{\mathcal{B}}$ with the equality corresponding to the case $\mathcal{B}=\mathcal{K}$ [31]. This is the minimization that we referred above as a functional minimization.
The Lanczos coefficients determine the matrix representation of the generator of evolution in Krylov space. Their specific role in quantum dynamics is still subject to investigation. The operator growth problem can be seen as a hopping single-particle in an one-dimensional semi-infinite chain with the Lanczos coefficients representing the hopping terms and the complexity in this case is the expectation value of the lattice index $n$ of the hopping particle; see, e.g., Ref. [12]. This interpretation is inspired by the fact that in the Krylov basis, the Hamiltonian has a tridiagonal form similar to the one of a tight-binding model, which describes the hopping of electrons in solids.

### v2

@@ Abstract
Investigating the time evolution of complexity in quantum systems entails evaluating the spreading of the system’s state across a defined basis in its corresponding Hilbert space. Recently, the Krylov basis has been identified as the one that minimizes this spreading. In this study, we develop a numerical exploration of the Krylov complexity in quantum states following a quench in the Lipkin-Meshkov-Glick model. Our results reveal that the long-term averaged Krylov complexity acts as an order parameter for this model. It effectively discriminates between the two dynamic phases induced by the quench, sharing a critical point with the conventional order parameter. Additionally, we examine the inverse participation ratio and the Shannon entropy in both the Krylov basis and the energy basis. A matching dynamic behavior is observed in both bases when the initial state possesses a specific symmetry. This behavior is analytically explained by establishing the equivalence between the Krylov basis and the pre-quench energy eigenbasis.
@@ I Introduction
Quantifying the complexity of physical phenomena has been a long-standing area of research [1]. Various measures, such as the entropy generated during the time evolution of a quantum state [2], algorithmic randomness [3], and quantum Kolmogorov complexity [4], have been utilized to characterize complexity. While existing measures capture specific features, some fail to encompass other aspects and exhibit ambiguities due to the specific basis choices involved in their definitions.
@@
Chaotic systems are expected to be more complex than integrable ones, a notion often attributed to phenomena like the butterfly effect  [5, 6]. In the realm of quantum systems, this effect is typically characterized by the exponential growth of the noncommutativity of local operators over time, as quantified by out-of-time order correlators. However, this exponential increase is not consistently observed in chaotic systems and may even manifest in nonchaotic ones [7, 8, 9].
@@
A novel measure of complexity has recently been proposed, relying on the spread of operators across a specifically ordered basis known as the Krylov basis [10]. This basis provides the description of the dynamics in terms of a single particle moving in a semi-infinite chain, with hopping rates set by the Lanczos coefficients obtained during the computation of the Krylov basis [11]. A large body of literature has followed, addressing fundamental properties of this measure [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23], its generalizations for other kinds of evolution [24, 25, 26, 27, 28, 29], and applications, e.g., to characterize phases of matter [30, 31, 32], and for quantum control [33, 34]. The Krylov complexity involves a choice of inner product between operators, raising questions about the optimal inner product for minimizing complexity. This challenge was addressed with the extension of the Krylov complexity for quantum state evolution presented in Ref. [35]. It was shown that the Krylov basis minimizes the proposed complexity definition. As a result, the Krylov complexity of quantum states, also known as the spread complexity, emerges as an unambiguous measure, eliminating the need for an inner product choice. It has been explored in the SYK model [10, 14], the XXZ model [16], the Su-Schrieffer-Heeger model [31], evolutions governed by dynamical symmetry groups associated with specific Lie algebras [17, 19, 26], and random matrix models [19, 35], among others. Additionally, it has been applied to analyze the transition from integrability to a chaotic regime [22, 36].
@@
The exploration of dynamics following a global quench in quantum many-body systems is widely used to probe a system in nonequilibrium statistical mechanics [37]. Global quenches induce rapid departures from equilibrium, along with coherence in the energy spectrum, producing intriguing effects, such as dynamical quantum phase transitions (DQPTs) [38, 39]. These transitions arise in quenched many-body systems and manifest as cusps in the Loschmidt echo or survival probability, i.e., non-analytic behavior in the time domain for certain initial states. This paper contributes to their study by investigating the evolution of Krylov complexity in quantum states following a quench in the paradigmatic Lipkin-Meshkov-Glick (LMG) model [40]. Our study of the LMG model, which may exhibit a DQPT depending on the quench [41], reveals that the long-time averaged Krylov complexity serves as an order parameter for the DQPT. The Krylov complexity manifests distinct behavior over time in both dynamical phases induced by the quench and remarkably mirrors the oscillations and revivals seen in one of the components of the total magnetization, a conventional dynamical order parameter in spin systems.
@@
In order to understand this fact, we conduct a numerical study of the inverse participation ratio (IPR) and the Shannon entropy over time in both the Krylov basis and the pre-quench energy basis. Remarkably, these two quantities can display identical behavior in both bases. Finally, by deriving the Krylov basis for the case of a quench starting at a null magnetic field, we analytically demonstrate its equivalence to the pre-quench energy basis. This analytical result explains the identical temporal behavior of IPR and Shannon entropy in these two bases and also why Krylov complexity is an order parameter for the considered model at null magnetic field.
Two early studies of the Krylov complexity in the LMG model have been reported. The first one [42] addresses scrambling dominated by classical saddle points in the model, and the second [32] identifies the spread complexity as a probe for the equilibrium quantum phase transition. However, none of these two works explores the phenomenon of DPQTs.
@@
The paper is structured as follows: Section II provides a review of key concepts related to dynamical phase transitions while Section III introduces the theory of Krylov complexity of quantum states. The goal of these two sections is twofold: setting the notation of the paper and making it self-contained. Section IV offers an overview of the dynamical phase transition in the LMG model, and Section V presents our results. We present our conclusions in Sec. VI.
@@ II Dynamical Quantum Phase Transition
also know as the survial amplitude [43, 44, 45, 46] and the Loschmidt amplitude [47].
The characterization of the DQPT relies on a formal analogy between $\mathcal{G}(t)$ and a special case of the boundary partition function $Z_{B}=\langle\psi_{1}|e^{-RH}|\psi_{2}\rangle$, with $R$ being the distance between the boundaries $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ [48].In this case, $R$ has to be interpreted as a spatiotemporal distance, in which one of its components is the complex time $it$ playing the role of the inverse temperature $\beta$, as guaranteed by relativistic quantum field theory. For more details, see also [49]. Thereby, formally, $\mathcal{G}(t)$ is the boundary partition function with $R=it$ and $|\psi_{1}\rangle=|\psi_{2}\rangle=|\psi_{0}\rangle$. The corresponding probability
@@
The DQPT is then defined in terms of the Fisher or Lee-Yang zeros [50, 51] of this partition function. Consider the return amplitude with complex time $t\to z=t+i\tau$. A transition occurs every time the zeros of $\mathcal{G}(z)$ crosses the real-time axis of the complex plane, and these crossings indicate the critical times of the dynamics. This analogy with the equilibrium quantum phase transition leads us to another important quantity in the dynamical case, a quantity that can be understood as the dynamical version of the free energy, the rate function
@@
with $N$ being the size of the system. At the critical times, $\mathcal{L}(t)$ is zero, and the rate function becomes nonanalytic. It is important to observe that this behavior only emerges in the thermodynamic limit where $N\rightarrow\infty$ and the Fisher zeros accumulate in a line or in a continuous region [38].
The above-described phenomenon was termed DPT-II to differentiate it from another related phenomenon known as DPT-I. DPT-II was initially investigated in the transverse field Ising model [47], and since then, it has been explored in various quantum many-body systems. These encompass short- and long-range interacting quantum spin chains [49, 52, 41, 53, 54, 55], two-dimensional spin systems [56, 57], nonintegrable systems [58], and optical systems [59], to name a few examples. For an in-depth review of DPT-II, covering both theoretical studies and experimental realizations, see Ref. [38].
@@
While DPT-II is investigated through the lens of the return probability, DPT-I is featured by the Landau order parameter of the system. This order parameter, a quantity whose derivative undergoes nonanalytic changes at a critical point of the quench parameter, serves to distinguish dynamical phases. Typically, the order parameter involves the time average of some physical quantity, such as the magnetization in spin chains. It is noteworthy that this dynamical critical point might not align with the critical point of the equilibrium QPT. DPT-I has been examined in various systems, and we point the reader to Refs. [39, 60] for comprehensive reviews on both its theoretical and experimental aspects.
@@
More recently, a connection between DPT-I and DPT-II and the excited-state quantum phase transition (ESQPT) in quantum many-body systems with infinite-range interactions has been established [62]. The authors in Ref. [62] define a generalized microcanonical ensemble by introducing three noncommuting charges and consider the presence of dynamical order parameters. In the thermodynamic limit, these order parameters are non-zero if the energy $E$ is less than the ESQPT critical energy $E_{c}$, but they always vanish when $E>E_{c}$. Concerning DPT-II, it was demonstrated that non-analyticities in the rate function $r(t)$ only occur if the energy of the system after the quench is greater than $E_{c}$, being absent if $E<E_{c}$. The same theoretical framework was applied to a finite-range interacting system in Ref. [63].
@@ III Krylov Complexity of Quantum States
Next, we introduce a notion of complexity of the quantum dynamics based on the spreading of $|\psi_{t}\rangle$ over the system Hilbert space, or rather, the subspace spanned by the quantum evolution known as the Krylov space. For this reason, this quantity is also referred to as spread complexity [35, 31].
@@
The successive application of $H$ to the initial state generates the set of quantum states
@@
which describes the spreading of the initial state over the Hilbert space during time evolution. The subspace spanned by $\{H^{n}|\psi_{0}\rangle\}_{n\geq 0}$ is known as the Krylov space. The measure of complexity that we consider in this work was introduced in Ref. [35], and it is given by
$C_{\mathcal{B}}(t)=\sum_{n}c_{n}|\langle\psi_{t}|B_{n}\rangle|^{2},$ (6)
where $\mathcal{B}$ represents, for the time being, a basis in the Hilbert space, whose elements are denoted as $\{|B_{n}\rangle\}_{n}$. Intuitively, we expect that complex dynamics lead to larger spreads over this basis. Thus, some constraints must be imposed on the coefficients $c_{n}$. For a real and positive-definite measure of complexity, regardless of the state of the system, $c_{n}>0$ is required. The coefficients should vary with $n$ and grow monotonically with $n$ to assign higher complexity to components $|\langle\psi_{t}|B_{n}\rangle|^{2}$ with higher index $n$. Note that the choice of constants coefficients $c_{n}=c$ is excluded as the complexity would reduce to the sum over all probabilities $p_{n}(t)=|\langle\psi_{t}|B_{n}\rangle|^{2}$ times $c$, yielding $c$ at all $t$.
A basis-independent complexity measure can be obtained from $C_{\mathcal{B}}(t)$ by performing a special minimization process, resulting in
@@
Of course, it is always possible to construct a basis at the initial time such that $|\psi_{0}\rangle=|B_{0}\rangle$ has a non-zero overlap only with one state of the basis, thereby minimizing complexity. Instead, we consider a functional minimization of (6) which takes into account the spread of the state. For this purpose, it is natural to look at the set $\{H^{n}|\psi_{0}\rangle\}_{n\geq 0}$ spanning the Krylov space since it includes only the portion of the Hilbert space visited by the system over time evolution.
The Krylov complexity, a precedent of the spread complexity, was first introduced in Ref. [10] in the context of operator evolution, i.e., in the Heisenberg picture. Based on this concept, a universal hypothesis for the maximum growth of local operators in quantum many-body systems was presented. Later, Balasubramanian and collaborators [35] extended the idea to the evolution of quantum states in the Schrödinger picture and proved that the basis which minimizes (6) for $c_{n}=n$ is the so-called Krylov basis. The Krylov basis is generated performing Gram-Schmidt orthogonalization on the set $\{H^{n}|\psi_{0}\rangle\}_{n\geq 0}$. We denote the Krylov basis as $\mathcal{K}$ and its elements as $\{|K_{n}\rangle\}_{n}$.
Alternatively, the Krylov basis can be generated using the Lanczos algorithm [11, 64], which is a well-known recursive method used to generate an orthogonal basis. Starting with the initial state as the first Krylov state $|K_{0}\rangle=|\psi_{0}\rangle$, the next state is obtained as $|K_{1}\rangle=\frac{1}{b_{1}}H|K_{0}\rangle$ with $b_{1}=\langle K_{1}|K_{1}\rangle^{1/2}$ being the normalization constant. The subsequent states $\{|K_{n}\rangle\}_{n\geq 2}$ are generated via the following recursion method
$|A_{n}\rangle=H|K_{n-1}\rangle-a_{n-1}|K_{n-1}\rangle-b_{n-1}|K_{n-2}\rangle,$ (8)
@@
Isolating the first term in the right-hand side of Eq. (8),
$H|K_{n-1}\rangle=a_{n-1}|K_{n-1}\rangle+b_{n}|K_{n}\rangle+b_{n-1}|K_{n-2}\rangle,$ (11)
we observe that the Hamiltonian is tridiagonal in the Krylov basis $\{|K_{n}\rangle\}_{n}$. We note that the Lanczos algorithm can lead to roundoff errors [65], affecting the orthogonality between the states. For this reason, it may be necessary to re-orthogonalize the states.
The authors of Ref. [35] proved that the Krylov basis minimizes Eq. (6) in a specific way. Formally, let
@@
We say that $S_{\mathcal{B}_{1}}<S_{\mathcal{B}_{2}}$ if there is some $k$ such that $C^{(m)}_{\mathcal{B}_{1}}=C^{(m)}_{\mathcal{B}_{2}}$ for $m<k$ and $C^{(m)}_{\mathcal{B}_{1}}<C^{(m)}_{\mathcal{B}_{2}}$ for $m=k$. Thus, for any basis $\mathcal{B}$, $S_{\mathcal{K}}\leq S_{\mathcal{B}}$ with the equality corresponding to the case $\mathcal{B}=\mathcal{K}$ [35]. This is the minimization that we referred to above as functional minimization.
The Lanczos coefficients determine the matrix representation of the generator of evolution in Krylov space. Their specific role in quantum dynamics is still subject to investigation. The complexity growth in both pictures can be seen as a hopping single particle in a one-dimensional semi-infinite chain with the Lanczos coefficients representing the hopping terms. Thus, it is expected that complex dynamics makes the wavepacket of the hopping particle delocalize quickly in the semi-infinite chain. It is also for this reason that one chooses $c_{n}=n$ in the definition of complexity, since upon this choice, the complexity is the average position of the hopping particle in the semi-infinite chain. In what follows, we refer to this chain as the Krylov chain.